### v1

@@ 5.2 Identifying hadronic PeVatrons
Currently there are two different concept of PeVatrons in the community, one
intended to address theoretically the first-principle question about the origin of CRs at the knee, and the other given from the observational perspective to describe sources showing signatures of PeV particles regardless of their contribution to the CR spectrum itself.
In the particle acceleration framework, the exact value of the particle maximum energy achieved depends on the spectral shape of the accelerated particles: e.g., in the case of a power-law spectrum with exponential cut-off, the maximum energy corresponds to the value where the flux deviates from the power law by 1/e, while in the case of a sharp cut-off it basically corresponds to the drop itself. Two important considerations are: first, that the maximum particle energy inferred today in a given system does not necessarily coincide with the maximum energy ever produced by it. While it does for stationary accelerators, as the wind termination shocks of stellar clusters, in impulsive accelerators as Supernova Remnants (SNRs) the escape of particles might deprive the shocks of the highest energy particles at an early stage. Our current in-situ observations can then only catch the still confined particle population (see Celli et al. (2019)). Hence, the correct assessment of PeVatron activity of a given system can only be achieved through dedicated models, and not simply by spectral fitting. Second, a further constraint for the CR PeVatron scenario concerns the slope of the injection spectrum, that is required to be between 2.2 and 2.4 to reproduce the local observations555Note that the injection spectrum does not necessarily correspond to the particle acceleration spectrum (see (Celli et al., 2019) for a detailed treatment concerning SNRs)..
The other definition, in turn, deals with the experimental signatures of PeVatron activity, e.g. the emission of UHE photons in the hadronic collisions of PeV protons and nuclei. On average, the ratio between the primary proton energy and the secondary gamma rays amounts to about a factor 10 across the spectral energy distribution, with the exception of the cut-off region, that in secondary particles appear milder than in primaries (Celli et al., 2020). As such, it is not sufficient to detect 100 TeV gamma rays to claim the presence of a PeVatron, rather the hadronic nature of the observed radiation should be probed and its spectral shape well sampled. To this extent, high-resolution spectral and angular observations are of paramount importance, and will be available with next-generation IACTs.
So far, LHAASO observations have revealed a few SNRs (Cao et al., 2023), characterised by featureless energy spectra and slopes significantly steeper than in the UHE band.
Nonetheless, in order to clearly assess the overall spectral trend of the emission and probe the physics of the source, it appears necessary to perform broadband fits, rater than being limited to the energy range of individual observations. In fact, identifying the presence of a cut-off or any other spectral variation would be crucial to the PeVatron science case: e.g. a cut-off might indicate either a drop in the available source power or the dominance of energy losses, depending on its shape, while an energy break might result from the occurrence of particle escape from the accelerator (Celli et al., 2019; Morlino & Celli, 2021). In summary, current measurements at the highest energies are still affected by large uncertainties to draw conclusive statements, requiring the collection of increased statistics and multi-wavelength analysis. In these regards, improved energy resolution will also be a key feature of future IACTs, possibly providing the resolving power to clearly identify spectral features at the highest energies.
@@
The list of LHAASO detected sources include many objects showing significant emission TeV, possibly competing to be hadronic PeVatrons. The only clear identification so far concerns the Crab Nebula, mostly dominated by leptonic emission. However, a hint for a spectral hardening at the highest energies probed by LHAASO suggests an additional component possibly be due to hadrons (Nie et al., 2022). The confirmation of this scenario awaits improving mapping of the emitting region at these extreme energies.
@@
It is hence of primary importance to investigate which source can be detected at 100 TeV by the next-generation IACTs. To do that, we extrapolate the flux reported from KM2A observations up to 100 TeV: in Fig. 7 we compare the resulting flux, , to the differential sensitivity of ASTRI and CTA. In particular, the left panel refers to observations at zenith angles, while the right panel concerns high-zenith angle () observations for CTA only. As it emerges from our estimation, CTA-South has the potential for investigating several objects at 100 TeV; among these, the source powering the region of 1LHAASO 1825-1337u and 1LHAASO J1809-1928u, possibly linked to pulsar emission (Cao et al., 2023), will be potentially visible already with a short 50-hour exposure. A deeper exposure coupled to high-zenith angle observations will enable all observatories to access a larger number of sources, despite their extensions.
@@ 5.3 Molecular clouds
Molecular clouds (MCs) illuminated by CRs constitute a class of Galactic gamma-ray emitters, typically detected in the GeV energy range (Aharonian et al., 2020; Peron et al., 2021; Baghmanyan et al., 2020). These are compact ( pc) regions of enhanced gas density ( cm) recognizable in the surveys of CO (J=21) line (Dame et al., 2001). When high-energy CRs impact the clouds, they produce (among other secondaries) gamma radiation proportional to both the density of the target gas and that of CRs, making MCs the most straightforward tracers of CRs far from the Earth.
@@
In the following, we evaluate the cloud visibility in the multi-TeV energy range. We consider the MCs listed in the catalog of Miville-Deschênes et al. (2017) as passive target for the local CR flux, and compare the expected gamma-ray flux emerging from hadronic collisions with the sensitivity of next-generation instruments, accounting for the cloud extension and the column density reported in the catalog. Following Peron & Aharonian (2022), we compute the resulting photon flux by assuming the locally measured CR flux (Orlando, 2018; Lipari & Vernetto, 2020), as:
@@
where is the cloud column density, is the angular area covered by the cloud, is the nuclear enhancement factor accounting for metals in both the CRs and the target gas, and is the inelastic proton-proton differential cross section (Peron & Aharonian, 2022). The assumption that the same CR spectrum across the Galaxy illuminates the clouds implies that the resulting hadronic gamma-ray flux only depends on their mass and distance. Indeed , in specific units. Following (Aharonian et al., 2020), we define as the ratio between the cloud mass (scaled to 10 solar masses) and the squared distance (in kpc units), and factorize the flux as , being the emissivity per single hydrogen atom. As a consequence, the visibility condition translates into a mere constraint regarding : defining the instrument sensitivity for given source extension as , then
@@
In Fig. 8 we compare the parameter characterizing the catalogued MCs to the ratio between sensitivity and emissivity at various energies and exposures. Also in this case, the zenith angle at which the cloud is observed might also influence the detection, especially at low energies. We consider here only the case of low-zenith angle observations (), and we indicate through the color code in Fig. 8 the zenith angle at which the clouds culminate at the Southern and Northern sites of CTA, the latter being located approximately at the same latitude as ASTRI and LHAASO. As it can be seen from the plot, the clouds with highest culminate at lower zenith angles in the Southern hemisphere, resulting in a possible detection of a few targets already with a 50-hour exposure by CTA-S. In Northern sites, on the other hand, the same clouds culminate at high-zenith angles, where observations are characterised by an increased energy threshold (see Appendix B). This is however limiting the prospects for the observation of the molecular clouds that are illuminated by the local CRs only, because a steep gamma-ray emissivity is expected in this case. Concerning LHAASO, only a few clouds are found to be possibly detectable with a 10-year long exposure. As a consequence, our result clearly suggests that passive MCs illuminated by the CR sea cannot be responsible for the unassociated sources in the first LHAASO catalog (Cao et al., 2023), which is based on a 933-day long exposure. However, future detection of MCs could enable to probe the Galactic CR spectrum up to TeV, which is a particularly interesting energy range because of the discrepant results currently shown by ground-based measurements (Aartsen et al., 2019).
@@
Nevertheless, one should consider that a large level of enhancement in the CR flux can be reached in the proximity of accelerators, due to the escaping particles (Aharonian & Atoyan, 1996; Gabici & Aharonian, 2007). E.g., the scenarios where Galactic clouds could be illuminated by SNR and YMSC escaped particles were studied in great details by Mitchell et al. (2021) and Celli et al. (2023), respectively. Furthermore, many among the clouds observed in gamma rays showed an enhancement as large as 5 compared to local (i.e. at Earth) cosmic-ray measured flux, even if not directly associated with an accelerator: such enhanced emissions from clouds were found mostly in the inner part of the Galaxy (Aharonian et al., 2020), but also in the local ( kpc) medium (Baghmanyan et al., 2020; Peron & Gabici, 2023). Similar enhancements would dramatically increase the detectability,allowing us to probe cloud spectra even at high energies, as done e.g. by the H.E.S.S. Collaboration (Sinha et al., 2021) for one of such systems, namely cloud 877 from the Rice et al. (2016) catalog. The extension of these type of observations to other objects, although challenging, would be of extreme relevance to probe the CR density elsewhere in the Galaxy especially at high energies, where a clear picture is not yet .
@@ 5.4 PWNe and pulsar halos
A new source class, named Pulsar/TeV halos, has emerged recently in the VHE surveys of the large FoV instrument HAWC. These sources are found in correspondence of pulsars, despite being much more extended than the associated surrounding PWN, therefore interpreted as a halo of particles escaping from the nebula. Interestingly, LHAASO observations have revealed that 40% of the first catalog sources are coincident with pulsars, possibly powerful enough to support UHE emission (de Oña Wilhelmi et al., 2022).
@@
Current IACT observations are limited by the extension of such halos, covering (and often going beyond) their entire FoV as in the case of Geminga (H. E. S. S. Collaboration et al. (2023)), representing a noticeable analysis challenge for the background estimation. Furthermore, significant flux differences have emerged in current measurements from IACTs and EASs of several sources because of the different integration region, making the case for observations with large FoV IACTs, like ASTRI, more compelling than ever.
@@
Interestingly, in order to explain the halo emission, a strong suppression of the diffusion coefficient is required, with strong implications regarding the amount of CR positron produced by these sources and in general on the propagation of Galactic CRs (López-Coto et al., 2022). However, Monogem and Geminga are found in the same region, therefore a local deformation of the magnetic field, which would impact the diffusion of particles regardless of the presence of the pulsars, cannot be excluded a priori. Despite this hypothesis being disfavoured theoretically (De La Torre Luque et al., 2022), further observations are necessary to conclude about the origin of the observed halos and to which extent they can influence the diffusion properties of the interstellar medium.
@@
At the moment, it is not clear whether these sources are the only pulsar halos in the sky or if any other exists, although some evidence for similar phenomena emerged in the most up-to-date catalogs of TeV sources. One candidate pulsar halo is the source HESS J1825-137, because its extension appears significantly larger than ordinary PWNe, suggesting that escape must be taking place from the nebula. Another candidate is LHAASO J0621+3755, one of the first sources detected in UHE gamma rays, its morphology being compatible with the scenario of escaped particles as claimed in Aharonian et al. (2021b). The observations with LHAASO-KM2A point towards a suppression of the diffusion coefficient to a similar level as measured in Geminga and Monogem. Again, high resolution spectral and morphological observations would be crucial for confirming the nature of this accelerator. Other candidate TeV halos have been announced by the HAWC collaboration, namely HAWC J0543+233 and HAWC J0635+070, both included in the second HAWC catalog (Abeysekara et al., 2017b): also in these cases, further morphological and spectral characterization are needed to be confirmed as halos, as current measurements could not resolve the radial profile of the emission neither the energy cut-off region with sufficient accuracy. All of these sources are expected to be powered by middle-age pulsars ( kyr). An exception is found in the pulsar of Vela, namely Vela X: Vela shows a dim extended emission detected both in radio and in the high-energy range, which is often interpreted as “relic PWN”, left behind after the interaction of the expanding PWN and the SNR reverse shock. The extension of the emission is however compatible also with a TeV halo scenario, making Vela X a possible intermediate case between the relic and the halo phase (Giacinti et al., 2020). Similarly, a TeV halo interpretation has been suggested to explain the extended emission detected by HAWC around the young pulsar PSR J0359+5414 (Albert et al., 2023). Furthermore, the aforementioned pulsar does not show a radio counterpart, analogously to the case of LHAASO J0621+3755. If this association is confirmed, it would imply that a larger variety of pulsars, compared to the present understanding, is able to produce halos.
@@
For these candidates, we provide in Fig. 9 a comparison between their differential fluxes and the calculated sensitivity of future IACTs and LHAASO for a 0.5 extension, compatibly with their reported sizes in TeVCat666http://tevcat.uchicago.edu and reported in the figure legend. The selected sources will be possibly visible with the benchmark exposures by the planned instruments, therefore allowing a comprehensive characterization of the emission around these pulsars, that so far is still missing. Investigating with high resolution the spatial profile of the flux of runaway particles might yield insight into radially-dependent diffusion properties. Such a goal is expected to be achieved in the future: in fact, the limited performance worsening when moving away from the axis center (Figs. 10, 11, 12, and 13) can be favourably exploited in such an application. We therefore expect high quality results in this field, although a more detailed evaluation needs to be performed to quantify the detectability of very large emitting regions as Geminga, e.g. accounting for the exact source geometry and spatial dependency of the spectrum.

### v2

@@ 5.2 Identifying hadronic PeVatrons
The definition of PeVatron is often considered in the CR and gamma-ray communities with different meanings: in the former, it is intended to address theoretically the first-principle question about the origin of CRs at the knee, while in the other it follows from an observational perspective to describe sources showing signatures of PeV particles, regardless of their contribution to the CR spectrum itself. Hence, when discussing about PeVatrons, it is necessary to clearly state which definition is concerned: in the present case, it will be based on the gamma-ray perspective. Both approaches are, however, model dependent: in fact, in realistic CR theories the maximum energy is often time-dependent, especially in impulsive accelerators as Supernova Remnants (SNRs) (Celli et al., 2019), such that detailed models are required to understand whether a given source has ever behaved as a PeVatron; in gamma-ray measurements, a caveat is that the maximum energy derived in fitting procedures strongly depends on the spectral shape of the cut-off (Celli et al., 2020), such that solid conclusions can only be derived as a result of data descriptions with regards to a given model, not simply by minimising the residuals with respect to a standard functional form.
With respect to the standard paradigm of Galactic CRs suggesting SNRs as main CR sources, we highlight that so far LHAASO observations have revealed a few UHE emitters (Cao et al., 2024a) in spatial coincidence with known SNRs, all sources being characterised by featureless and steep gamma-ray energy spectra. While the collection of increased statistics and multi-wavelength analysis are necessary to clearly assess the overall spectral trend of their emission and to probe the presence of spectral variations, it is unlikely for these objects to behave as main CR contributors.
@@
In these regards, the improved energy resolution will be a key feature of future IACTs, possibly providing the resolving power to distinctly identify spectral features at the highest energies. The list of LHAASO detected sources include many other objects showing significant emission TeV. Among these, the only clear identification so far concerns the Crab Nebula, mostly dominated by leptonic emission, although a hint for a spectral hardening at the highest energies probed by LHAASO suggests an additional component possibly be due to hadrons (Nie et al., 2022). The confirmation of this scenario awaits improving mapping of the emitting region at these extreme energies.
@@
Given the primary importance to investigate which of the LHAASO detected sources can be observed at 100 TeV by the next-generation IACTs, both in the context of exploring spectral features and associate unidentified ones, we now focus on the KM2A detected sources from the first LHAASO catalogue. As no information is available in the catalogue concerning the maximum photon energy measured for each of them, we extrapolate the reported fluxes up to 100 TeV. In Fig. 7 we compare the resulting flux at 100 TeV, , to the differential sensitivity of ASTRI and CTA at the very same energy: in particular, the left panel refers to observations at zenith angles, while the right panel concerns high-zenith angle () observations for CTA only. As it emerges from our estimation, CTA-South has the potential for investigating several objects at 100 TeV; among these, the source powering the region of 1LHAASO 1825-1337u and 1LHAASO J1809-1928u, possibly linked to pulsar emission (Cao et al., 2024a), will be potentially visible already with a short 50-hour exposure. A deeper exposure coupled to high-zenith angle observations will enable all observatories to access a larger number of sources, despite their extensions. All detectable sources are labelled with their name in Fig. 7.
@@ 5.3 Molecular clouds
Molecular clouds (MCs) illuminated by CRs constitute a class of Galactic gamma-ray emitters, typically detected in the GeV energy range (Aharonian et al., 2020; Peron et al., 2021; Baghmanyan et al., 2020). These are compact ( pc) regions of enhanced gas density ( cm-3) recognizable in the surveys of CO (J=21) line (Dame et al., 2001). When high-energy CRs impact the clouds, they produce (among other secondaries) gamma radiation proportional to both the density of the target gas and that of CRs, making MCs the most straightforward tracers of CRs far from the Earth.
@@
In the following, we evaluate the cloud visibility of the MCs listed in the catalogue of Miville-Deschênes et al. (2017) for the next-generation instruments in the multi-TeV energy range, by comparing the expected gamma-ray flux emerging from hadronic collisions of the local CR flux with their sensitivity, accounting for the cloud extension and the column density reported in the catalogue. Following Peron & Aharonian (2022), we compute the resulting photon flux by assuming the locally measured CR flux (Orlando, 2018; Lipari & Vernetto, 2020).
@@
The assumption that the same CR spectrum across the Galaxy illuminates the clouds implies that the resulting hadronic gamma-ray flux only depends on their mass and distance. Indeed , in specific units. Following (Aharonian et al., 2020), we define as the ratio between the cloud mass (normalised to 105 solar masses) and the squared distance (in kpc units), and factorise the expected gamma-ray flux as , being the emissivity per single hydrogen atom. As a consequence, the visibility condition translates into a mere constraint regarding : defining the instrument sensitivity for given source extension as , then
@@
In Fig. 8 we compare the parameter characterising the catalogued MCs to the ratio between sensitivity and emissivity at various energies and for different exposures. Also in this case, the zenith angle at which the cloud is observed might influence the detection, especially at low energies. We consider here only the case of low-zenith angle observations (), and we indicate through the colour code of each cloud the zenith angle of their culmination at the southern and northern sites of CTA, the latter being located approximately at the same latitude as ASTRI and LHAASO. Potentially detectable clouds are highlighted with a label reporting their unique identifier as provided in the Miville-Deschênes et al. (2017) catalogue. To this extent, we selected clouds with an expected gamma-ray flux overcoming the detection threshold for extended sources with at least one of the considered IACTs and culminating at zenith angles lower than 60∘ at the corresponding site. As it can be seen from the plot, the clouds with highest culminate at lower zenith angles in the southern hemisphere, resulting in a possible detection of a few targets already with a 50-hour exposure by CTA-S. In northern sites, on the other hand, the same clouds culminate at high-zenith angles, where observations are characterised by an increased energy threshold (see Appendix C). This is however limiting the prospects for the observation of the MCs that are illuminated by local CRs only, because a steep gamma-ray emissivity is expected in this case. Concerning LHAASO, a few clouds are found to be possibly detectable only with a 10-year long exposure. As a consequence, our result clearly suggests that passive MCs illuminated by the CR sea cannot be responsible for the unassociated sources in the first LHAASO catalogue (Cao et al., 2024a), which is based on a 933-day long exposure. However, future detection of MCs could enable the Galactic CR spectrum to be probed up to TeV, which is a particularly interesting energy range because of the discrepant results currently shown by ground-based measurements (Aartsen et al., 2019).
@@
For a few of these clouds, namely cloud 57 and 78 the Fermi-LAT analysis (Peron et al., 2021) unveiled a spectrum similar to the locally measured one. Nevertheless, one should consider that a large level of enhancement in the CR flux is expected in the proximity of accelerators, due to the escaping particles (Aharonian & Atoyan, 1996; Gabici & Aharonian, 2007). E.g., the scenarios where Galactic clouds could be illuminated by SNR and YMSC escaped particles were studied in great details by Mitchell et al. (2021) and Celli et al. (2023), respectively. Furthermore, many among the clouds observed in gamma rays showed an enhancement as large as 5 compared to local (i.e. at Earth) CR measured flux, even if not directly associated with an accelerator: the increased emissions from clouds were found mostly in the inner part of the Galaxy (Aharonian et al., 2020), but also in the local ( kpc) medium (Baghmanyan et al., 2020; Peron & Gabici, 2023). Similar enhancements would dramatically increase the detection prospects, allowing us to probe cloud spectra even at high energies. One compelling case is the case of cloud 17, which coincides with the cloud 877 of the Rice et al. (2016). Its enhanced emission reported by Fermi-LAT (Aharonian et al., 2020) stimulated follow-up observations with H.E.S.S. (Sinha et al., 2021), that reported a clear detection and proposed a dedicated analysis method for such low-surface brightness sources. The extension of these type of observations to other objects, although challenging, would be of extreme relevance to probe the CR density elsewhere in the Galaxy especially at high energies, where a clear picture is not yet there.
@@ 5.4 PWNe and pulsar halos
A new source class, named pulsar/TeV halos, has emerged recently in the VHE surveys of the large FoV instrument HAWC. These sources are found in correspondence of pulsars, despite being much more extended than the associated surrounding PWN, therefore interpreted as a halo of particles escaping from the nebula. Interestingly, LHAASO observations have revealed that 40% of the first catalogue sources are coincident with pulsars, possibly powerful enough to support UHE emission (de Oña Wilhelmi et al., 2022).
@@
Current IACT observations are limited by the extension of such halos, covering (and often going beyond) their entire FoV as in the case of Geminga (H. E. S. S. Collaboration et al. (2023a)), representing a noticeable analysis challenge for the background estimation. Furthermore, significant flux differences have emerged in current measurements from IACTs and EASs of several sources because of the different integration region, making the case for observations with large FoV IACTs, such as ASTRI, more compelling than ever.
@@
Interestingly, in order to explain the halo emission, a strong suppression of the diffusion coefficient or a transition to a different propagation region (Recchia et al., 2021), is required, with implications regarding the amount of CR positron produced by these sources and in general on the propagation of Galactic CRs (López-Coto et al., 2022). However, Monogem and Geminga are found in the same region, therefore a local deformation of the magnetic field, which would impact the diffusion of particles regardless of the presence of the pulsars, cannot be excluded a priori. Despite this hypothesis being disfavoured theoretically (De La Torre Luque et al., 2022), further observations are necessary to conclude about the origin of the observed halos and to which extent they can influence the diffusion properties of the interstellar medium.
@@
At the moment, it is not clear whether these sources are the only pulsar halos in the sky or if any other exists, although some evidence for similar phenomena emerged in the most up-to-date catalogues of TeV sources. One candidate pulsar halo is the source HESS J1825-137, because its extension appears significantly larger than ordinary PWNe, suggesting that escape must be taking place from the nebula. Another candidate is LHAASO J0621+3755, one of the first sources detected in UHE gamma rays, its morphology being compatible with the scenario of escaped particles as claimed in Aharonian et al. (2021c). The observations with LHAASO-KM2A point towards a suppression of the diffusion coefficient to a similar level as measured in Geminga and Monogem. Again, high resolution spectral and morphological observations would be crucial for confirming the nature of this accelerator. Other candidate TeV halos have been announced by the HAWC Collaboration, namely HAWC J0543+233 and HAWC J0635+070, both included in the second HAWC catalogue (Abeysekara et al., 2017b): also in these cases, further morphological and spectral characterisation are needed to be confirmed as halos, as current measurements could not resolve the radial profile of the emission neither the energy cut-off region with sufficient accuracy. We note that energetic emission is measured by LHAASO from a direction consistent with both these sources, namely LHAASO J0543+2311u and LHAASO J0635+0619, both observed by KM2A the former even at UHE. All of the aforementioned sources are expected to be powered by middle-age pulsars ( kyr). An exception is found in the pulsar of Vela, namely Vela X: Vela shows a dim extended emission detected both in radio and in the high-energy range, which is often interpreted as ’relic PWN’, left behind after the interaction of the expanding PWN and the SNR reverse shock. The extension of the emission is however compatible also with a TeV halo scenario, making Vela X a possible intermediate case between the relic and the halo phase (Giacinti et al., 2020). Similarly, a TeV halo interpretation has been suggested to explain the extended emission detected by HAWC around the young pulsar PSR J0359+5414 (Albert et al., 2023). Such a pulsar, also detected by Fermi, does not show a radio counterpart, analogously to the case of LHAASO J0621+3755. If this association is confirmed, it would imply that a larger variety of pulsars, compared to the present understanding, is able to produce halos. The spatial association with LHAASO J0359+5406 remains to be confirmed, given that also PSR B0355+54 could in principle power its emission.
@@
For these candidates, we provide in Fig. 9 a comparison between their differential fluxes and the calculated sensitivity of future IACTs and LHAASO for a 0.5∘ extension, compatibly with their reported sizes in TeVCat666http://tevcat.uchicago.edu and reported in the figure legend. The selected sources will be possibly visible with the benchmark exposures by the planned instruments, therefore allowing a comprehensive characterisation of the emission around these pulsars, that so far is still missing. Investigating with high resolution the spatial profile of the flux of runaway particles might yield insight into radially dependent diffusion properties. Such a goal is expected to be achieved in the future: in fact, the limited performance worsening when moving away from the axis centre (Figs. 13, 14, 15, and 16) can be favourably exploited in such an application. We therefore expect high quality results in this field, although a more detailed evaluation needs to be performed to quantify the detectability of very large emitting regions as Geminga, e.g. accounting for the exact source geometry and spatial dependency of the spectrum.
6 Discussion
In the light of defining the best candidates for future IACT observations, we provide in Fig. 10 gamma-ray source fluxes at 50 TeV from both the LHAASO-KM2A and the H.G.P.S. catalogues, as compared to sensitivities of IACTs.
The choice of 50 TeV as a reference energy follows from the measurements by LHAASO-KM2A sources. On the other hand, for H.E.S.S. sources we obtain the expected 50 TeV fluxes by extrapolating the measured spectra. Given the impact of the zenith angle on IACT sensitivity, we show source gamma-ray fluxes as a function of the minimum observation angle in the northern and southern sites, respectively in the upper and lower panels of the figure. We expect that the major observational efforts will be most-likely devoted, in the next decade, to unveil the nature of unidentified sources and measure possible cut-off features to find out any source that does not show a cut-off earlier than PeV. To facilitate the identification of these objects, we explicitly indicate in the figure, next to source names, their class (unidentified, SNR, PWN, binaries, etc.) as well as their best-fit spectral shape (PL, where a cut-off is not detected), whenever this information is available.
We can hence define a list of the most promising targets for each hemisphere, selecting the top five brightest unidentified sources characterised by featureless spectra, observable at zenith angles below 60∘ in the corresponding sites.
For the northern sites, we highlight:
HESS J1908+063/1LHAASO J1908+0615u: these two spatially overlapping sources have been revealed by different surveys, but it remains to be clarified if they are the same or not. We note that the flux measured by LHAASO-KM2A at 50 TeV is much lower than the extrapolated flux from H.E.S.S., therefore if those are indeed the same source, a spectral break would be necessary to connect the two spectra. Multiple counterparts are possible PWN and SNR but also the Juchert 3 star cluster (Aharonian et al., 2009; Celli et al., 2024)
1LHAASO J1814-1636u: this UHE source is only detected by KM2A with angular extension of . Two SNRs are located nearby, namely G14.1-00.1 ( away) and SNR G014.3+00.1 ( away). However, both SNR candidates have radio shell size, hence inconsistent with that of gamma-ray emission. An unidentified GeV source, 4FGL J1816.2-1654c, is also found away from the position of the TeV emission, but so far no clear counterpart emerged.
HESS J1825-137, 1LHAASO J1825-1337u, and 1LHAASO J1825-1418 (good targets also for the southern observatories): the H.E.S.S. source is spatially overlapping with two LHAASO sources, though the association is unclear. Notably in this case, the source observed by H.E.S.S. has a flux significantly smaller than the LHAASO sources, such that it again appears necessary to determine whether the emission can be ascribed to the same source or if multiple emitters are present. In both cases, an emerging component at 50 TeV is particularly interesting. The H.E.S.S. sources, identified as PWN by the collaboration, is one of the candidate TeV halos that we discussed in Sec. 5.
HESS J1809-193/1LHAASO J1809-1918u: analogously to the previous case, also here, two apparently coincident sources have an inconsistent spectrum, where the LHAASO measured flux is much larger than the one extrapolated from H.E.S.S; in this case, though, the counterpart remains obscure. In the past, this source was thought to be a PWN, but other possible counterparts such as SNRs and molecular later emerged, making the recognition of the dominating emission mechanism not straightforward. (H. E. S. S. Collaboration et al., 2023b).
HESS J1843-033/1LHAASO J1843-0335u: little is known about this source, that seems to be characterised by two components; in particular, the H.E.S.S. source coincides with UBC 353, a weak and quite old star cluster from the Gaia catalogue (Celli et al., 2024).
For the southern sites, in turn, we suggest in-depth observations to be carried for:
HESS J1702-420: this unidentified source is a candidate PeVatron, due to its hard energy spectrum up to the highest energies probed (Abdalla et al., 2021), such that it is among the most interesting sources to investigate.
HESS J1616-508: this source is also unidentified in the H.E.S.S. catalogue and, due to the conspicuous nature of the diffuse gas in there, it is possibly composed of a yet-undetected central accelerator producing a particle flux interacting with the gas (Lau et al., 2017).
HESS J1458-608: with regards to this unidentified source, its VHE gamma-ray emission is thought to be associated with the pulsar PSR J1459-6053 such that it seems likely to be an undetected PWN, on the basis of a spatial coincidence with an energetic pulsar and the absence of other plausible multi-wavelength counterparts (Devin et al., 2021).
HESS J1708-443: this is a faint extended source that is possibly associated to a PWN, although a SNR is also a possible counterpart (H. E. S. S. Collaboration et al., 2011).